Give an example of a subring of a ring, say A, that is not an ideal of A


1
Expert's answer
2021-12-06T17:16:11-0500

Consider the ring "(\R,+,\cdot)" of real numbers. Let "A=\mathbb Q" be the set of rational numbers. Since for "a,b\in\mathbb Q" we get that "a-b\in\mathbb Q" and "a\cdot b\in\mathbb Q," we conclude that "(\mathbb Q,+,\cdot)" is a subring of the ring "(\R,+,\cdot)." On the other hand, for "\sqrt{2}\in\R" and "1\in\mathbb Q" we get that "\sqrt{2}\cdot 1=\sqrt{2}\notin\mathbb Q," and hence "(\mathbb Q,+,\cdot)" is not an ideal of the ring "(\R,+,\cdot)."